# Robustness-Aware 3D Object Detection in Autonomous Driving: A Review and Outlook

### v1

@@ III-A Monocular 3D object detection
Monocular 3D object detection refers to performing 3D object detection using only one camera which aims to infer the 3D position, size, and orientation of target objects from a single image [274]. In recent years, monocular 3D object detection has gained increasing attention due to its advantages of low cost, low power consumption, and ease of deployment in real-world applications. However, monocular methods face many challenges, owing to the insufficient 3D information in monocular pictures, such as accurately localizing the 3D position, handling occluded scenes, etc. Overcoming these challenges relies on leveraging depth information to supplement the missing 3D information in monocular images. Typically, most approaches employ depth estimation tasks to acquire depth information from images. However, monocular depth estimation is an ill-posed and highly challenging task, prompting researchers to dedicate significant efforts to optimize the accuracy and stability of depth estimation.
@@ III-A1 Prior-guided monocular 3D object detection
In recent years, prior-guided monocular methods [291, 292, 293, 294, 295, 296, 297, 17, 49, 298, 299, 15, 300, 5, 43, 16, 68, 6, 301, 18, 39, 47, 20, 296] have continuously explored how to utilize the hidden prior knowledge of object shapes and scene geometry in images to address the challenges of monocular 3D object detection. This effective integration of prior knowledge is crucial for mitigating the uncertainty and ill-posed nature of monocular 3D object detection problems. By introducing pre-trained subnetworks or auxiliary tasks, prior knowledge can provide additional information or constraints to assist in the accurate localization of 3D objects and enhance detection precision and robustness.
Widely adopted prior knowledge in 3D objects includes object shapes [17, 338, 339, 296, 297, 293], geometric consistency [340, 300, 43, 5, 43, 16, 6], temporal constraints [47, 302], and segmentation information [296, 20]. Object shape provides insights into the appearance and structure of the object, aiding in more accurate inference of the spatial position and pose of the object. Geometric consistency knowledge assists the model in better understanding the relative positional relationships between objects in the scene, thereby improving detection consistency and robustness. Temporal constraints consider the continuity and stability of the object across different frames, providing vital clues for object detection. Additionally, leveraging segmentation information enables the model to better comprehend semantic information in the images, facilitating precise localization and identification of objects.
As a result, current works are dedicated to further exploring and utilizing prior knowledge to enhance the performance and robustness of monocular 3D object detection by integrating prior knowledge with deep learning approaches, thus driving continuous development and innovation in this field. The early algorithm Mono3D [50] first assumes the 3D object is on a fixed ground plane, and then uses the prior 3D shape of the vehicle to reconstruct the bounding box in 3D space. In subsequent work, Deep MANTA [292] uses keypoints and 3D CAD models to predict 3D objects. Pose-RCNN [341] learns viewpoint-specific subclass information from 3D CAD models to capture shape, viewpoint information, and potential occlusion patterns of objects. MonoPSR [49] generates 3D candidate frames for each object in the scene by using the fundamental relationships of the pinhole camera model and a well-established 2D object detector.
With a deeper understanding and application of prior knowledge, it is believed that significant progress will be achieved in monocular 3D object detection in the future, bringing breakthroughs and opportunities to the fields of computer vision and intelligent systems.
@@ III-A2 Camera-only monocular 3D object detection
Camera-only monocular 3D object detection [21, 15, 302, 66, 57, 39, 299, 65, 67, 48, 5, 6, 53, 52, 303, 7, 304, 305, 301, 68] is a method that utilizes images captured by a single camera to detect and localize 3D objects. Camera-only monocular methods employs convolutional neural networks (CNNs) to directly regress 3D bounding box parameters from the images, enabling the estimation of the spatial dimensions and poses of objects in three dimensions. Drawing inspiration from the architectural design of 2D detection networks, this direct regression method can be trained in an end-to-end manner, facilitating holistic learning and inference for 3D objects. The unique challenge of monocular 3D object detection lies in inferring the 3D position, dimensions, and orientation of objects solely from a single image, without relying on additional depth maps or point cloud data. Consequently, the direct regression approach demonstrates practicality and broad applicability. By learning features from the images, convolutional neural networks can predict the 3D information of the objects. Through end-to-end training, the network gradually optimizes its parameters to enhance the accurate extraction of 3D information. This direct regression method streamlines the entire detection process and reduces reliance on supplementary information, thereby improving the algorithm’s robustness and generalization capability. Nevertheless, monocular 3D object detection still presents challenges, such as occlusion, viewpoint variations, and changes in lighting conditions, which may impact the accuracy of 3D detection. The representative work Smoke[21] abandons the regression of 2D bounding boxes and predicts the 3D box for each detected target by combining the estimation of individual keypoints with the regression of 3D variables.
@@ III-A3 Depth-assisted monocular 3D object detection
Depth estimation plays a crucial role in depth-assisted monocular 3D object detection. To achieve more accurate monocular detection results, numerous studies [306, 10, 69, 37, 3, 4, 307, 45, 36, 62] leverage pre-trained auxiliary depth estimation networks. Specifically, the process begins by transforming monocular images into depth images using pre-trained depth estimators, such as MonoDepth[342]. Subsequently, two primary methodologies are employed to handle depth images and monocular images.
Remarkable progress has been made in Pseudo-LiDAR detectors that use a pre-trained depth estimation network to generate Pseudo-LiDAR representations[60, 306, 61, 62]. However, there is a huge performance gap between Pseudo-LiDAR and LiDAR-based detectors because of the errors in image-to-LiDAR generation. Thus, Hong et al.[59] attempted to transfer deeper structural information from point clouds to assist monocular image detection. By leveraging the mean-teacher framework, they aligned the outputs of the LiDAR-based teacher model and the camera-based student model at both the feature-level and response-level, aiming to achieve cross-modal knowledge transfer. Such a depth-assisted monocular 3D object detection, by effectively integrating depth information, not only enhances detection accuracy but also extends the applicability of monocular vision to tasks involving 3D scene understanding.
@@ III-B Stereo-based 3D object detection
Stereo-based 3D object detection is designed to identify and localize 3D objects using a pair of stereo images. Leveraging the inherent capability of stereo cameras to capture dual perspectives, stereo-based methods excel in acquiring highly accurate depth information through stereo matching and calibration, a feature that sets them apart from monocular camera setups. Despite these advantages, stereo-based methods still face a considerable performance gap when compared to LiDAR-based counterparts. Furthermore, the realm of 3D object detection from stereo images remains relatively underexplored, with only limited research endeavors dedicated to this domain. Specifically, these approaches involve the utilization of image pairs captured from distinct viewpoints to estimate the 3D spatial depth of each object.
III-B1 2D-detection based methods
@@
Traditional 2D object detection frameworks can be modified to address stereo detection problems. Stereo R-CNN [23] employs an image-based 2D detector to predict 2D proposals, generating left and right regions of interest (RoIs) for the corresponding left and right images. Subsequently, in the second stage, it directly estimates the parameters of 3D objects based on the previously generated RoIs. This paradigm has been widely adopted by follow works [13, 308, 309, 310, 311, 312, 24, 313].
III-B2 Pseudo-LiDAR based methods
@@
The disparity map predicted from stereo images can be transformed into a depth map and further converted into a pseudo-LiDAR points. Consequently, similar to monocular detection methods, pseudo-LiDAR representations can also be employed in stereo-based 3D object detection approaches. These methods aim to enhance the disparity estimation in stereo matching to achieve more accurate depth predictions. Regarding the contribution of depth in 3D detection, Wang et al. [3] is a pioneer in introducing the Pseudo-LiDAR representation. This representation is generated by an image with a depth map, requiring the model to perform a depth estimation task to assist in detection. Subsequent work has followed this paradigm and made optimizations by introducing additional color information to augment pseudo point cloud [45], auxiliary tasks(instance segmentation [36], foreground and background segmentation [343] and domain adaptation [40]) and coordinate transform scheme [8, 306]. It is worth noting the insightful work proposed by Ma et al. called PatchNet [306]. Specifically, the authors challenges the conventional idea of leveraging the pseudo-LiDAR representation for monocular 3D object detection. By encoding 3D coordinates for each pixel, PatchNet can attain a comparable monocular detection result without pseudo-LiDAR representation. This observation indicates that the power of the pseudo LiDAR representation stems from the coordinate transformation rather than the point cloud representation itself.
@@ III-C Multi-view 3D object detection
Recently, multi-view 3D object detection has demonstrated superior accuracy and robustness compared to the aforementioned monocular and stereo 3D object detection approaches. In contrast to LiDAR-based 3D object detection, the latest panoramic Bird’s Eye View (BEV) approaches eliminate the need for high-precision maps, elevating the detection from 2D to 3D. This advancement has led to significant developments in multi-view 3D object detection. In comparison to previous reviews [1, 344, 345, 346, 347, 242, 2, 348, 349, 350, 269, 252, 270, 271, 272, 273, 274, 275], there has been extensive research on effectively leveraging multi-view images for 3D object detection. In multi-camera 3D object detection, a key challenge lies in recognizing the same object across different images and aggregating object features from multiple view inputs. The current approach involves uniformly mapping multi-view to the Bird’s Eye View (BEV) space, which is a common practice. Therefore, multi-view 3D object detection, also referred to as BEV-camera-only 3D object detection, revolves around the core challenge of unifying 2D views into the BEV space. Based on different spatial transformations, this can be categorized into two main methods: one approach is depth-based methods [25, 315, 316, 92, 93, 95, 96, 98, 100, 317, 318, 319, 320, 321, 322, 323, 324], represented by the LSS [316], also known as the 2D to 3D transformation. The other approach is query-based methods [99, 87, 97, 26, 89, 90, 94, 166, 319, 325, 326, 327, 328, 329, 330, 331, 332, 333, 334, 91, 335, 336, 337] represented by the DETR3D [88], achieving a query from 3D to 2D.
@@ III-C1 Depth-based Multi-view methods
The direct transformation from 2D to BEV space poses a significant challenge. As shown in Fig. 3, LSS [316] was the first to propose a depth-based method, utilizing 3D space as an intermediary. This approach involves initially predicting the grid depth distribution of 2D features and then elevating these features to voxel space. This method holds promise for more effectively achieving the transformation from 2D to BEV space. Following LSS [316], CaDDN [52] adopted a similar depth representation approach. It employed a network structure akin to LSS, primarily for predicting categorical depth distribution. By compressing voxel-space features to BEV space, it performed the final 3D detection. It is worth noting that CaDDN is not part of multi-view 3D object detection, but rather single-view 3D object detection, which has influenced subsequent research on depth. The main distinction between LSS [316] and CaDDN [52] lies in CaDDN’s use of actual ground truth depth values to supervise its prediction of categorical depth distribution, resulting in an outstanding depth network capable of more accurately extracting 3D information from 2D space. This line of research has sparked a series of subsequent studies, such as BEVDet [315], its temporal version BEVDet4D [95], and BEVDepth [25]. These studies are of great significance in advancing the transformation from 2D to 3D space and enabling more accurate object detection in the BEV space, providing valuable insights and directions for the relevant field’s development. Furthermore, some studies have addressed the issue of insufficient depth solely by encoding height information. These studies have found that with increasing distance, the depth disparity between the car and the ground rapidly diminishes [317, 100].
@@ III-C2 Query-based Multi-view methods
Under the influence of Transformer [351, 352, 353, 354] technology, query-based Multi-view methods retrieve 2D spatial features from 3D space. Inspired by Tesla’s perception system, DETR3D[88] introduces 3D object queries to address the aggregation of multi-view features. It accomplishes this by clipping image features from different perspectives and projecting them into 2D space using learned 3D reference points, thus obtaining image features in the Bird’s Eye View (BEV) space. Query-based Multi-view methods, contrary to Depth-based Multi-view methods, acquire sparse BEV features by employing a reverse querying technique, fundamentally impacting subsequent query-based developments [99, 87, 97, 26, 89, 90, 94, 166, 319, 325, 326, 327, 328, 329, 330, 331, 332, 333, 334, 91, 335, 336, 337]. However, due to the potential inaccuracies associated with explicit 3D reference points, PETR [89], as shown in Fig. 4, influenced by DETR [355] and DETR3D [88], adopts an implicit positional encoding method for constructing the BEV space, influencing subsequent works[90, 336]. Meanwhile, some methods [225, 356] do not explicitly construct Bird’s Eye View (BEV) features.
@@ III-D Analysis: Accuracy, Latency, Robustness
Currently, the 3D object detection solutions based on Bird’s Eye View (BEV) perception are rapidly advancing. Despite the existence of numerous reviews[1, 344, 345, 346, 347, 242, 2, 348, 349, 350, 269, 252, 270, 271, 272, 273, 274, 275], a comprehensive review of this field remains inadequate. It is noteworthy that Shanghai AI Lab and SenseTime Research have provided a thorough review [357] of the technical roadmap for BEV solutions. However, unlike existing reviews [1, 344, 345, 346, 347, 242, 2, 348, 349, 350, 269, 252, 270, 271, 272, 273, 274, 275], which primarily focus on the technical roadmap and the current state of the art, we considers crucial aspects such as autonomous driving safety perception. Following an analysis of the technical roadmap and the current state of development for camera-based solutions, we intend to base our discussion on the foundational principles of ‘Accuracy, Latency, Robustness.’ We will integrate the perspectives of safety perception to guide the practical implementation of safety perception in autonomous driving.

### v2

@@ III-A Monocular 3D object detection
Monocular 3D object detection refers to performing 3D object detection using only one camera, which aims to infer the 3D positions, sizes, and orientations of objects from a single image [131]. In recent years, monocular 3D object detection has gained increasing attention due to its advantages of low cost, low power consumption, and ease of deployment in real-world applications. However, monocular methods face many challenges, owing to the insufficient 3D information in monocular pictures, such as accurately localizing 3D positions, handling occluded scenes, and so on. Overcoming these challenges relies on leveraging depth information to supplement the missing 3D information in monocular images. Typically, most approaches employ depth estimation tasks to acquire depth information from images. However, monocular depth estimation is an ill-posed and highly challenging task, prompting researchers to dedicate significant efforts to optimizing the accuracy and stability of depth estimation.
@@ III-A1 Prior-guided monocular 3D object detection
In recent years, prior-guided monocular methods [352, 162, 163, 164, 165, 166, 167, 21, 168, 169, 11, 170, 7, 19, 12, 26, 8, 184, 13, 18] have continuously explored how to utilize the hidden prior knowledge of object shapes and scene geometry in images to address the challenges of monocular 3D object detection. The effective integration of this prior knowledge is crucial for mitigating the uncertainty and ill-posed nature of inherent in monocular 3D object detection problems. By introducing pre-trained subnetworks or auxiliary tasks, prior knowledge can provide additional information or constraints to assist in the accurate localization of 3D objects and enhance detection precision and robustness.
Widely adopted prior knowledge in 3D objects includes object shapes [353, 354, 355, 166, 167, 163], geometric consistency [356, 170, 19, 7, 19, 12, 8], temporal constraints [357, 180], and segmentation information [166]. Object shape provides insights into the appearance and structure of an object, aiding in more accurate inference of the spatial position and pose of the object. Geometric consistency knowledge assists the model in better understanding the relative positional relationships between objects in the scene, thereby improving detection consistency and robustness. Temporal constraints consider the continuity and stability of an object across different frames, providing vital clues for object detection. Additionally, leveraging segmentation information enables the model to better comprehend semantic information in images, facilitating precise localization and identification of objects. As a result, current works are dedicated to further exploring and utilizing prior knowledge to enhance the performance and robustness of monocular 3D object detection by integrating prior knowledge with deep learning approaches, thus driving continuous development and innovation in this field.
@@ III-A2 Camera-only monocular 3D object detection
Camera-only monocular 3D object detection [179, 11, 180, 18, 169, 24, 25, 7, 8, 22, 181, 9, 182, 183, 184, 26] is a kind of methods that utilize images captured by a single camera to detect and localize 3D objects. Camera-only monocular methods employ convolutional neural networks (CNNs) to directly regress 3D bounding box parameters from images, enabling the estimation of the spatial dimensions and poses of objects in three dimensions. Inspired by 2D detection networks, this direct regression methods can be trained end-to-end, facilitating holistic learning and inference for 3D objects. The unique challenge of monocular 3D object detection lies in inferring objects’ 3D positions, dimensions, and orientations solely from a single image without relying on additional depth maps or point cloud data. Consequently, the direct regression approaches demonstrate practicality and broad applicability. By learning features from images, CNNs can predict the 3D information of objects. The network gradually optimizes its parameters through end-to-end training to enhance the accurate extraction of 3D information. These direct regression methods streamline the entire detection process and reduce the reliance on supplementary information, improving the algorithms’ robustness and generalization capability. Nevertheless, monocular 3D object detection still presents challenges, such as occlusion, viewpoint variations, and lighting conditions, which may affect the accuracy of 3D detection. The representative work Smoke[179] abandons the regression of 2D bounding boxes and predicts the 3D box for each detected object by combining the estimation of individual key points with the regression of 3D variables.
@@ III-A3 Depth-assisted monocular 3D object detection
Depth estimation plays a crucial role in depth-assisted monocular 3D object detection. To achieve more accurate monocular detection results, numerous studies [201, 27, 202, 20] leverage pre-trained auxiliary depth estimation networks. Specifically, the process begins by transforming monocular images into depth images using pre-trained depth estimators, such as MonoDepth[358]. Subsequently, two primary methodologies are employed to handle depth images and monocular images. Remarkable progress has been made in Pseudo-LiDAR detectors that use a pre-trained depth estimation network to generate Pseudo-LiDAR representations[359, 201]. However, there is a significant performance gap between Pseudo-LiDAR and LiDAR-only detectors due to the errors in image-to-LiDAR generation. Thus, Hong et al.[360] attempted to transfer deeper structural information from point clouds to assist monocular image detection. By leveraging the mean-teacher framework, they aligned the outputs of the LiDAR-only teacher model and the Camera-only student model at both the feature-level and the response-level, aiming to achieve cross-modal knowledge transfer. Such depth-assisted monocular 3D object detection, by effectively integrating depth information, not only enhances detection accuracy but also extends the applicability of monocular vision to tasks involving 3D scene understanding.
@@ III-B Stereo-based 3D object detection
Stereo-based 3D object detection is designed to identify and localize 3D objects using a pair of stereo images. Leveraging the inherent capability of stereo cameras to capture dual perspectives, stereo-based methods excel in acquiring highly accurate depth information through stereo matching and calibration. This is a feature that distinguishes them from monocular camera setups. Despite these advantages, stereo-based methods still face a considerable performance gap when compared to LiDAR-only counterparts. Furthermore, the realm of 3D object detection from stereo images remains relatively underexplored, with only limited research efforts dedicated to this domain. Specifically, these approaches involve the utilization of image pairs captured from distinct viewpoints to estimate the 3D spatial depth of each object.
III-B1 2D-detection-based methods
@@
Traditional 2D object detection frameworks can be modified to address stereo detection problems. Stereo R-CNNs [285] employ an image-based 2D detector to predict 2D proposals, generating left and right regions of interest (RoIs) for the corresponding left and right images. Subsequently, in the second stage, they directly estimate the parameters of 3D objects based on the previously generated RoIs. This paradigm has been widely adopted by subsequent works [214, 215, 216, 195, 217, 14, 218].
III-B2 Pseudo-LiDAR-only methods
@@
The disparity map predicted from stereo images can be transformed into a depth map and further converted into pseudo-LiDAR points. Consequently, similar to monocular detection methods, pseudo-LiDAR representations can also be employed in stereo-based 3D object detection approaches. These methods aim to enhance disparity estimation in stereo matching to achieve more accurate depth predictions. Regarding the contribution of depth in 3D detection, Wang et al. [6] are pioneers in introducing the Pseudo-LiDAR representation. This representation is generated by using an image with a depth map, requiring the model to perform a depth estimation task to assist in detection. Subsequent works have followed this paradigm and made optimizations by introducing additional color information to augment pseudo point cloud [20], auxiliary tasks (instance segmentation [361], foreground and background segmentation [362] and domain adaptation [363]) and coordinate transform scheme [364, 201]. To achieve both high accuracy and high responsiveness, Meng et al. [365] propose a lightweight Pseudo-LiDAR 3D detection system. These studies indicate that the power of the pseudo LiDAR representation stems from the coordinate transformation rather than the point cloud representation itself.
III-B3 Volume-based methods
The general procedure of volume-based methods is to generate a cost volume from the left and the right images to represent disparity information, which is then utilized in the subsequent detection process. Volume-based methods bypass the pseudo-LiDAR representation and perform 3D object detection directly on 3D stereo volumes. These methods have inherited the traditional matching idea, but most computations now rely on 3D convolutional networks, such as those found in references [235, 366, 245, 242, 254, 17, 238, 236]. For example, the pioneering work GC-Net[235] uses an end-to-end neural network for stereo matching, obviating the need for any post-processing steps, and regressively computes disparity from a cost volume constructed by a pair of stereo features. GwcNet[366] employs a proposed group-wise correlation method to construct the cost volume. LEAStereo[245] utilizes NAS technology to select the optimal structure for the 3D cost volume. GANet[242] designs a semi-global aggregation layer and a local guidance aggregation layer to further improve accuracy. ACVNet[254] introduces an attention concatenation unit to generate more accurate similarity metrics. DSGN [17] proposes a 3D geometric volume derived from stereo matching networks and applies a grid-based 3D detector on the volume for 3D object detection. LIGA-Stereo[238] uses a LiDAR-based detector as a teacher model to guide geometry-aware feature learning. ESGN[236] achieves efficient stereo matching through the efficient geometry-aware feature generation (EGFG) module. Due to the benefits of large-scale training data and end-to-end training, deep learning-based stereo methods have achieved outstanding results[243].
@@ III-C Multi-view 3D object detection
Recently, multi-view 3D object detection has demonstrated superior accuracy and robustness compared to monocular and stereo 3D object detection approaches. In contrast to LiDAR-only 3D object detection, the latest panoramic Bird’s Eye View (BEV) approaches eliminate the need for high-precision maps, elevating the detection from 2D to 3D. This advancement has led to significant developments in multi-view 3D object detection. In comparison to previous reviews [4, 126, 123, 130, 131], there has been extensive research on effectively leveraging multi-view images for 3D object detection. A key challenge in multi-camera 3D object detection is recognizing the same object across different images and aggregating object features from multiple view inputs. The current approach, a common practice, involves uniformly mapping multi-view to the Bird’s Eye View (BEV) space. Therefore, multi-view 3D object detection, also called BEV-camera-only 3D object detection, revolves around the core challenge of unifying 2D views into the BEV space. Based on different spatial transformations, this can be categorized into two main methods. Ones are depth-based methods [15, 255, 256, 33, 35, 257, 258, 268, 259, 263, 367, 368, 264], represented by the LSS [256], also known as 2D to 3D transformation. The others are query-based methods [267, 34, 16, 31, 32, 84, 268, 269, 369, 270, 271, 272, 273, 274, 275, 276, 277, 278, 279, 280], represented by DETR3D [30], making a query from 3D to 2D.
@@ III-C1 Depth-based Multi-view methods
The direct transformation from 2D to BEV space poses a significant challenge. LSS [256] was the first to propose a depth-based method, utilizing 3D space as an intermediary. This approach involves initially predicting the grid depth distribution of 2D features and then elevating these features to voxel space. This method holds promise for achieving the transformation from 2D to BEV space more effectively. Following LSS [256], CaDDN [22] adopted a similar depth representation approach. It employed a network structure akin to LSS, primarily for predicting categorical depth distribution. By compressing voxel-space features into BEV space, it performed the final 3D detection. It is worth noting that CaDDN is not part of multi-view 3D object detection but rather single-view 3D object detection, which has influenced subsequent research on depth. The main distinction between LSS [256] and CaDDN [22] lies in CaDDN’s use of actual ground truth depth values to supervise its prediction of categorical depth distribution, resulting in a superior depth network capable of more accurately extracting 3D information from 2D space. This line of research has sparked a series of subsequent studies, such as BEVDet [255], its temporal version BEVDet4D [33], and BEVDepth [15]. These studies are significant in advancing the transformation from 2D to 3D space and enabling more accurate object detection in the BEV space, providing valuable insights and directions for the field’s development. Furthermore, some studies have addressed the issue of insufficient depth solely by encoding height information. These studies have found that with increasing distance, the depth disparity between the car and the ground rapidly diminishes [257, 35].
@@ III-C2 Query-based Multi-view methods
Under the influence of Transformer technology, such as in the works [370, 371, 372, 373], query-based Multi-view methods retrieve 2D spatial features from 3D space. Inspired by Tesla’s perception system, DETR3D[30] introduces 3D object queries to address the aggregation of multi-view features. It achieves this by extracting image features from different perspectives and projecting them into 2D space using learned 3D reference points, thus obtaining image features in the Bird’s Eye View (BEV) space. Query-based Multi-view methods, as opposed to Depth-based Multi-view methods, acquire sparse BEV features by employing a reverse querying technique, fundamentally impacting subsequent query-based developments [267, 34, 16, 31, 32, 84, 268, 269, 369, 270, 271, 272, 273, 274, 275, 276, 277, 278, 279, 280]. However, due to the potential inaccuracies associated with explicit 3D reference points, PETR [31], influenced by DETR [374] and DETR3D [30], adopts an implicit positional encoding method for constructing the BEV space, influencing subsequent works[32, 279].
@@ III-D Analysis: Accuracy, Latency, Robustness
Currently, the 3D object detection solutions based on Bird’s Eye View (BEV) perception are rapidly advancing. Despite the existence of numerous reviews[4, 126, 123, 130, 131], a comprehensive review of this field remains inadequate. It is noteworthy that Shanghai AI Lab and SenseTime Research have provided a thorough review [375] of the technical roadmap for BEV solutions. However, unlike existing reviews [4, 126, 123, 130, 131], which primarily focus on the technical roadmap and the current state of the art, we consider crucial aspects such as autonomous driving safety perception. Following an analysis of the technical roadmap and the current state of development for Camera-only solutions, we intend to base our discussion on the foundational principles of ‘Accuracy, Latency, and Robustness’. We will integrate the perspectives of safety perception to guide the practical implementation of safety perception in autonomous driving.